the score of shs students in their first quarter in statistic and probability brainly were analyzed and found to have a mean of 54.5and standard deviation of 7.5. find a 90% confidence interval


1
Expert's answer
2022-04-27T14:25:27-0400

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(54.5-1.6449\times\dfrac{7.5}{\sqrt{n}}, 54.5+1.6449\times\dfrac{7.5}{\sqrt{n}})"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "54.5-1.6449\times\dfrac{7.5}{\sqrt{n}}<\mu< 54.5+1.6449\times\dfrac{7.5}{\sqrt{n}},"

which indicates that we are 90% confident that the true population mean "\mu"

is contained by the interval 


"(54.5-1.6449\times\dfrac{7.5}{\sqrt{n}}, 54.5+1.6449\times\dfrac{7.5}{\sqrt{n}})."